OpenStudy (priyar):

A thin tube sealed at both ends is 100 cm long . It lies horizontally,the middle 20cm containing mercury and 2 equal ends containing air at std. atm. pressur. If the tube is now turned to a vertical position, by what amount will the mercury be displaced? (given cross scetion is uniform)

OpenStudy (priyar):

@ParthKohli @ganeshie8

Parth (parthkohli):

oh god why do you hate me so much :(

Parth (parthkohli):

hint: boyle's law

OpenStudy (samigupta8):

Lolxcc @pk

OpenStudy (samigupta8):

Okk @priyar Did u get it ??

OpenStudy (priyar):

no...

Parth (parthkohli):

hey I'm going out for a walk but\[P\ell = P_1 \ell_1 = P_2 \ell_2\]\[P_2 = P_1 + \rho_{mercury}\cdot 10 ~\rm cm \cdot g\]

OpenStudy (samigupta8):

Better write as l only because we can assume pressure equal to h...

OpenStudy (samigupta8):

I mean l2=l1+10

OpenStudy (priyar):

why 10?

OpenStudy (priyar):

@samigupta8?

OpenStudy (priyar):

why h=10?

OpenStudy (priyar):

pls explain properly..

OpenStudy (priyar):

why is h=10??? why not h=20? anyone?

OpenStudy (samigupta8):

Sorry i wrote it wrong ...it should be what u told just now 20

OpenStudy (samigupta8):

I just saw parth's post and thought that the mercury column was 10cm long..

OpenStudy (priyar):

how did you both make the same mistake?!!

OpenStudy (samigupta8):

I didn't see the value in ur question for mercury column length

OpenStudy (priyar):

and..density of mercury is not given!

OpenStudy (samigupta8):

It doesn't matter ...bcoz the pressure are being taken in term of length so we can assume that l2=l1+20

OpenStudy (samigupta8):

Sorry @priyar ..i m wrong ... I were considering a very diff case from ur question.. Wait lemme correct it in my notebook and then i should tell you how to move with this problem

OpenStudy (priyar):

hmm..okay.. thats why i got confused..

OpenStudy (samigupta8):

What are the options pls..?

Parth (parthkohli):

sorry guys I just came back that 10 cm was a typo

OpenStudy (priyar):

ok.. but density of mercury isn't given..

Parth (parthkohli):

that's ok... actually the thing is that pressure is measured as length of mercury too. 1 atm = 76 cm mercury.

OpenStudy (priyar):

i just came to ask if we were to use that..

Parth (parthkohli):

yes, we have to use that. in the upper column of gas, the pressure is the atmospheric pressure (76 cm). and in the lower one, it is atmospheric pressure + 20 cm = 96 cm mercury.

OpenStudy (samigupta8):

U don't need it even... U can do it in this fashion... P2=P1+rho *20*g Then P2= 40*76/80-h And P1=40*P/h And now write P as 76 cm of Hg Now eq iz this 40*76/80-h -40*76/h=20

OpenStudy (samigupta8):

Here i assumed p2 is pressure in upper column of the tube And p1 in lower column

OpenStudy (priyar):

how P2= 40*76/80-h?

OpenStudy (samigupta8):

Sorry i wrote it wrong p2 is pressure in lower column

OpenStudy (priyar):

still explain how you got P2..

OpenStudy (samigupta8):

P*40=p1*h P*40=p2(80-h)

OpenStudy (priyar):

isn't that (60 - h) ?.. and the answer is not matching.....

OpenStudy (samigupta8):

Why would it be 60-h?

OpenStudy (priyar):

coz 40+20 = 60..right.. and 2 is at 60cm from the top.. @samigupta8

OpenStudy (priyar):

@IrishBoy123 @imqwerty

OpenStudy (priyar):

@ParthKohli @Michele_Laino

OpenStudy (vincent-lyon.fr):

The answer should be a displacement of 5.18 cm.

OpenStudy (priyar):

@Vincent-Lyon.Fr that's correct! pls explain how to solve..

OpenStudy (vincent-lyon.fr):

Since section is constant, PV = cst is the same as \(P.l=cst\) If pressures are expressed in cm Hg, and lengths in cm, then \(\rho_{Hg}.g=1\) It makes equations to solve easier when expressed numerically. You have to solve: \(P_o.l_o=P_1.l_1=P_2.l_2\) (Boyle's law) \(P_o=76\) and \(l_o=40\) \(l_1+l_2=80\) \(P_2-P_1=20\)

OpenStudy (priyar):

how ρHg.g=1 ?

OpenStudy (priyar):

pressure = ρHg.g.h right?

OpenStudy (priyar):

@Vincent-Lyon.Fr

OpenStudy (vincent-lyon.fr):

If h = 1 cm, then P = \(\rho_{Hg}.g.h=1\; cm Hg\)

OpenStudy (priyar):

yes...okay

OpenStudy (priyar):

are these what you have taken as l1, l2